# Disorder-induced proximate quantum spin ice phase in $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$

### v1

@@ III.3 Bulk characterization
As shown in Fig. 4(a,b), a.c. susceptibility measurements on single-crystal $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$  with zero d.c. field reveal a gradual slowing down of spin dynamics below 1 K. This is manifested in the frequency-dependent behavior of the real part, $\chi^{\prime}$, whose peak shifts to higher temperatures with increasing frequency. A complete spin-freezing transition occurs at $T_{f}\approx 0.15$ K, where $\chi^{\prime}$ drops nearly to zero. By applying a damped-harmonic oscillator model Topping and Blundell (2019); Ortiz et al. (2024) to $\chi^{\prime}$, we estimate that $\tau_{\mathrm{slow}}$ ranges from $10^{-6}$ s to $10^{-2}$ s below 1 K [Fig. 4(i), Fig. A3]. The frequency dependence of the freezing temperature $T_{f}$, determined from the position of the cusp in the imaginary part $\chi^{\prime\prime}$ shown in Fig. 4(b), follows Arrhenius behavior with activation energy $E_{a}=2.7(7)$ K [Fig. 4(c)], consistent with previous results on powder samples Ortiz et al. (2024).
The zero-field heat capacity $C_{p}$ is shown in Fig. 4(d) and exhibits a pronounced nuclear Schottky anomaly that approaches the theoretical maximum value of $\sim 7~\mathrm{J}/(\mathrm{mol}_{\mathrm{Pr^{3+}}}\cdot\mathrm{K})$, corresponding to fully frozen Pr moments Kimura et al. (2013); Tang et al. (2022) at base temperature. The recovered zero-point entropy, obtained by subtracting the hyperfine contribution (see Appendix B), approaches the full Pauling value near 10 K. Taken together, these results suggest that while the system enters a fully spin-frozen state at base temperature, a quantum spin-ice (QSI) phase may emerge above the spin-freezing transition, characterized by a “slow” timescale of $\tau_{\mathrm{slow}}\sim 10^{-5}$ s Tomasello et al. (2019).
@@ III.4 Neutron spectroscopy
At $T=0.5$ K and 10 K, we observe quasi-elastic scattering extending up to $\hbar\omega\sim 0.6$ meV, indicative of persistent spin fluctuations above $T_{f}$. To characterize the energy dependence of this quasi-elastic scattering, we adopt a generalized Lorentzian model Böni and Shirane (1986),
@@
where the first Gaussian term accounts for elastic contributions including nuclear scattering, static magnetic correlations, and residual background (see Fig. 5(i,ii,iii)). The Gaussian width $\sigma\approx 0.0087$ meV is fixed by the instrumental resolution. The second term describes the quasi-elastic magnetic scattering. The phenomenological exponent $\epsilon(\omega)$ is introduced to capture the high-energy tails, defined as $\epsilon(\omega)=1$ for $|\omega|\leq\Gamma$ and $\epsilon(\omega)=1+\alpha(|\omega|-\Gamma)/\Gamma$ for $|\omega|>\Gamma$ Böni and Shirane (1986). The best-fit results are shown as dashed lines in Fig. 5(b,c), yielding characteristic energy widths of $\Gamma=0.062(3)$ meV at $T=0.5$ K and $\Gamma=0.10(1)$ meV at $T=10$ K. These correspond to timescales $\tau_{\mathrm{fast}}\sim h/\Gamma$ of approximately 0.07 ns and 0.04 ns, respectively. We interpret these timescales as signatures of the “fast” spin dynamics predicted for QSI systems Tomasello et al. (2019). In both fits, small values of the exponent $\alpha$ in $\epsilon(\omega)$ were obtained: $\alpha=0.13(2)$ at $T=0.5$ K and $\alpha=0.05(13)$ at $T=10$ K.
@@
At $T=0.5$ K, we observe a characteristic “pinch-point” pattern in the elastic channel, integrated over $\hbar\omega\in[-0.2,0.2]$ meV [Fig. 5(d)]. Upon cooling into the spin-frozen state at $T=0.05$ K, the $\bm{Q}$-dependence of the elastic scattering changes dramatically, as shown in Fig. 5(e). The resulting pattern is reminiscent of the magnetic diffuse scattering observed in CSI materials with strong dipolar interactions such as $\text{Ho}_{2}\text{Ti}_{2}\text{O}_{7}$Bramwell et al. (2001); Kanada et al. (2002) and $\text{Dy}_{2}\text{Ti}_{2}\text{O}_{7}$Yavors’kii et al. (2008); Samarakoon et al. (2022), but deviates even more from the ideal pinch-point structure. Instead, the scattering becomes concentrated near reciprocal space positions associated with a $\bm{q}=(100)$ propagation vector, suggesting a tendency towards incipient $\bm{q}=(100)$ order as the system undergoes freezing. The $\bm{Q}$-dependence of the gapped excitation observed at $T=0.05$ K [Fig. 5(a)] is revealed in Fig. 5(g), showing a starfish-shaped distribution reminiscent of a smeared pinch-point pattern, similar to inelastic magnetic features previously observed in several QSI candidates Petit et al. (2016a, b); Wen et al. (2017); Sibille et al. (2018). A similar $\bm{Q}$-dependence is observed for the quasielastic scattering at $T=0.5$ K over the same energy range $\hbar\omega\in[0.2,0.6]$ meV [Fig. 5(f)], although the features appear more diffuse. For higher energy transfers $\hbar\omega\geq 0.6$ meV [Fig. 5(h, j)], no significant $\bm{Q}$-dependence is observed. Additional data are presented in Figs. A7,  A9, and A10.
@@ IV Discussion
The second mechanism assumes that the spin-freezing transition is driven by the spatial variance of the transverse field, $\delta h$, exceeding a critical threshold $\delta h_{c}$ Savary and Balents (2017); Benton (2018). In this regime, certain ring-exchange terms become significantly enhanced relative to others, favoring the condensation of vison excitations. We interpret the observed suppression of quasielastic scattering below $T_{f}$ as a signature of vison condensation, which suppresses the low-energy quantum fluctuations characteristic of a quantum spin ice (QSI). Theoretically, vison condensation has also been predicted to induce $(100)$ magnetic order Chen (2016), which indeed appears incipient in the frozen state [Fig. 5(e)].
@@
Our results mark an important step toward realizing a U(1) QSL state, as they indicate that $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$ lies in close proximity to the ideal U(1) QSL in two key respects. First, in the temperature range above $T_{f}$ but below the onset of spin-ice correlations at $\sim$1 K, $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$ exhibits hallmarks of a QSL state: the characteristic wavevector dependence of the diffuse scattering Benton et al. (2012) and the presence of two well-separated magnetic timescales Tomasello et al. (2019). Second, the spin-frozen state that develops below $T_{f}$ appears to emerge from the QSL via a confinement transition, in contrast to floating-zone–grown $\text{Pr}_{2}\text{Zr}_{2}\text{O}_{7}$ samples Kimura et al. (2013); Petit et al. (2016a); Wen et al. (2017), which remain in a paramagnetic regime separated from the QSI phase by a first-order transition.
@@ Appendix A Model for the gapped excitation in spin-frozen state
As discussed in Section IV, the spin-frozen phase below $T_{f}\approx 0.15$ K hosts a gapped excitation at $\hbar\omega=0.23$ meV with a starfish-like $\bm{Q}$ dependence, reminiscent of inelastic features reported in other QSI candidates Petit et al. (2016a, b); Wen et al. (2017); Sibille et al. (2018). In $\text{Pr}_{2}\text{Zr}_{2}\text{O}_{7}$, analogous excitations were attributed to CEF transitions between singlets of a split non-Kramers doublet, with the $\bm{Q}$ dependence shaped by finite nearest-neighbor correlations $J_{1}$ Wen et al. (2017); Benton (2018). Within a mean-field framework Petit et al. (2016a); Martin et al. (2017), the elastic and inelastic spectral weights scale as $|\cos\theta_{i}|^{2}$ and $|\sin\theta_{i}|^{2}$, respectively, where $\theta_{i}$ is the angle between the pseudospin (defined in Eq. 1) and the local $\hat{\bm{z}}$ axis. Our finding that the inelastic channel carries only $\sim$40% of the total weight indicates a moderate deviation of the pseudospins from the Ising axis and suggests that the system retains features of the paramagnetic phase in the phase diagram [Fig. 1].
To characterize further the behavior of these CEF excitations, we follow Ref. Martin et al. (2017) and introduce a mean-field model in which disorder in Ising pyrochlores with a non-Kramers ground-state doublet is represented by a site-dependent transverse field $h_{i}$ acting on each Pr3+ ion Savary and Balents (2017); Benton (2018). The system is described by the Hamiltonian given by Eq. 4.
When $h_{i}$ is finite, the pseudo-spin $\sigma_{i}$ at site $i$ is no longer aligned strictly along the local $\langle 111\rangle$ $\hat{\bm{z}}$ direction. Following Refs. Petit et al. (2016a); Martin et al. (2017), we consider a local mean-field Hamiltonian for site $i$,
@@
where $a_{1}$, $b_{1}$, $a_{2}$, and $b_{2}$ are normalized coefficients, satisfying $|a_{1}|^{2}+|b_{1}|^{2}=1$ and $|a_{2}|^{2}+|b_{2}|^{2}=1$. Compared to the convention adopted in Ref. Petit et al. (2016a), where a local spherical angle $\theta_{i}$ is defined between the pseudospin and the $\hat{\bm{z}}$ direction, the parameters used here are related by: $a_{1}\simeq\sin(\theta_{i}/2)$, $b_{1}\simeq\cos(\theta_{i}/2)$, $a_{2}\simeq\cos(\theta_{i}/2)$, and $b_{2}\simeq-\sin(\theta_{i}/2)$.
Since neutrons primarily couple to the $\hat{\bm{z}}-$component of the real spin for systems with non-Kramers ground state doublet Chen (2017), and the projected spin operator $J_{z}$ within the non-Kramers doublet is proportional to the pseudo-spin operator $\sigma_{z}$ Petit et al. (2016a), the neutron scattering intensities in the elastic and inelastic channels can be expressed as:
@@ Appendix B Subtraction of nuclear Schottky anomaly in heat capacity
We analyzed the nuclear Schottky anomaly in $C_{p}$ using two models: (1) the analytical model of Kimura et al. Kimura et al. (2013), with $\mu_{\mathrm{hyp}}^{\mathrm{Pr}}=1.9~\mathrm{\upmu_{B}}$ and no scaling factor, as in Ref. Kimura et al. (2013); Ortiz et al. (2024); and (2) a statistical model treating the system as a multilevel scheme with evenly spaced energy levels, using a level spacing of $\epsilon_{r}=0.011$ meV. Both models successfully reproduce the broad maximum in $C_{p}$ near 0.1 K. We note that our heat capacity data and the parameters obtained from model (1) are very similar to those of $\text{Pr}_{2}\text{Zr}_{2}\text{O}_{7}$ sample 1 reported in Ref. Tang et al. (2022). The temperature dependence of the magnetic entropy $S_{m}$, shown in Fig. 4(ii), was determined by subtracting both the Gopal-model nuclear contribution and the lattice contribution (estimated from nonmagnetic La2Sn2O7) from the measured $C_{p}$, and integrating the resulting magnetic contribution $C_{m}/T$. We emphasize that subtraction using either the Kimura or the Gopal model yields essentially identical results for the magnetic entropy.
@@ Appendix C Additional information for $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$ samples
Magnetization versus field measurements at $T=2$ K on single-crystal $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$  samples with masses of approximately 5 mg are shown in Fig. A4(c). The induced magnetic moments $M$ at the maximum applied field $\mu_{0}H=7$ T are measured to be $1.64~\mu_{\mathrm{B}}/\mathrm{Pr^{3+}}$, $1.33~\mu_{\mathrm{B}}/\mathrm{Pr^{3+}}$, and $1.19~\mu_{\mathrm{B}}/\mathrm{Pr^{3+}}$ for fields applied along the $[100]$, $[111]$, and $[110]$ directions, respectively. For comparison, the expected saturated moments in a pyrochlore lattice with Ising anisotropy along the local $\langle 111\rangle$ directions and effective ferromagnetic (FM) coupling between nearest neighbors are $g_{J}J/\sqrt{3}$ for the $[100]$ direction, $g_{J}J/2$ for $[111]$, and $g_{J}J/\sqrt{6}$ for $[110]$ Fukazawa et al. (2002). Both the observed ordering of the induced moments (from largest to smallest) and their relative ratios are consistent with the predictions for a magnetic pyrochlore with easy-axis FM anisotropy. Quantitative deviations may arise from slight misalignments during sample mounting, mass uncertainty due to residual impurity phases, or the fact that full saturation may not have been achieved at $\mu_{0}H=7$ T. The temperature dependence of the dc magnetic susceptibility was measured with the magnetic field applied along the $[100]$, $[111]$, and $[110]$ crystallographic directions using the same single-crystal samples. To obtain a representative bulk behavior, we compute a powder-averaged dc susceptibility for a cubic system using the expression $\chi_{\mathrm{powder}}^{\mathrm{DC}}=(3\chi_{100}^{\mathrm{DC}}+6\chi_{110}^{\mathrm{DC}}+4\chi_{111}^{\mathrm{DC}})/13$ Bonville et al. (2016), and plot the results in Fig. A4(d,e). This powder-averaged dc susceptibility is also used to normalize the real part of the ac susceptibility $\chi^{\prime}$ presented in Fig. 5(a). A Curie–Weiss fit to the high-temperature regime yields a Curie–Weiss temperature of $\theta_{\mathrm{CW}}=-0.23(2)$ K and an effective moment of $\mu_{\mathrm{eff}}=2.78(1)~\mu_{\mathrm{B}}/\mathrm{Pr^{3+}}$, in good agreement with the powder measurements reported in Ref. Ortiz et al. (2024).
To extract the “slow” timescale of spin dynamics Tomasello et al. (2019) from the real part $\chi^{\prime}$ of the ac susceptibility, we employ a damped harmonic oscillator model Topping and Blundell (2019); Ortiz et al. (2024):

### v2

@@ III.3 Bulk characterization
As shown in Fig. 4(a,b), a.c. susceptibility measurements on single-crystal $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$  with zero d.c. field reveal a gradual slowing down of spin dynamics below 1 K. This is manifested in the frequency-dependent behavior of the real part, $\chi^{\prime}$, whose peak shifts to higher temperatures with increasing frequency. A complete spin-freezing transition occurs at $T_{f}\approx 0.15$ K, where $\chi^{\prime}$ drops nearly to zero. By applying a damped-harmonic oscillator model [42, 35] to $\chi^{\prime}$, we estimate that $\tau_{\mathrm{slow}}$ ranges from $10^{-6}$ s to $10^{-2}$ s below 1 K [Fig. 4(i), Fig. A3]. The frequency dependence of the freezing temperature $T_{f}$, determined from the position of the cusp in the imaginary part $\chi^{\prime\prime}$ shown in Fig. 4(b), follows Arrhenius behavior with activation energy $E_{a}=2.7(7)$ K [Fig. 4(c)], consistent with previous results on powder samples [35].
The zero-field heat capacity $C_{p}$ is shown in Fig. 4(d) and exhibits a pronounced nuclear Schottky anomaly that approaches the theoretical maximum value of $\sim 7~\mathrm{J}/(\mathrm{mol}_{\mathrm{Pr^{3+}}}\cdot\mathrm{K})$, corresponding to fully frozen Pr moments [21, 44] at base temperature. The recovered zero-point entropy, obtained by subtracting the hyperfine contribution (see Appendix B), approaches the full Pauling value near 10 K. Taken together, these results suggest that while the system enters a fully spin-frozen state at base temperature, a quantum spin-ice (QSI) phase may emerge above the spin-freezing transition, characterized by a “slow” timescale of $\tau_{\mathrm{slow}}\sim 10^{-5}$ s [36].
@@ III.4 Neutron spectroscopy
At $T=0.5$ K and 10 K, we observe quasi-elastic scattering extending up to $\hbar\omega\sim 0.6$ meV, indicative of persistent spin fluctuations above $T_{f}$. To characterize the energy dependence of this quasi-elastic scattering, we adopt a generalized Lorentzian model [45],
@@
where the first Gaussian term accounts for elastic contributions including nuclear scattering, static magnetic correlations, and residual background (see Fig. 5(i,ii,iii)). The Gaussian width $\sigma\approx 0.0087$ meV is fixed by the instrumental resolution. The second term describes the quasi-elastic magnetic scattering. The phenomenological exponent $\epsilon(\omega)$ is introduced to capture the high-energy tails, defined as $\epsilon(\omega)=1$ for $|\omega|\leq\Gamma$ and $\epsilon(\omega)=1+\alpha(|\omega|-\Gamma)/\Gamma$ for $|\omega|>\Gamma$ [45]. The best-fit results are shown as dashed lines in Fig. 5(b,c), yielding characteristic energy widths of $\Gamma=0.062(3)$ meV at $T=0.5$ K and $\Gamma=0.10(1)$ meV at $T=10$ K. These correspond to timescales $\tau_{\mathrm{fast}}\sim h/\Gamma$ of approximately 0.07 ns and 0.04 ns, respectively. We interpret these timescales as signatures of the “fast” spin dynamics predicted for QSI systems [36]. In both fits, small values of the exponent $\alpha$ in $\epsilon(\omega)$ were obtained: $\alpha=0.13(2)$ at $T=0.5$ K and $\alpha=0.05(13)$ at $T=10$ K.
@@
At $T=0.5$ K, we observe a characteristic “pinch-point” pattern in the elastic channel, integrated over $\hbar\omega\in[-0.2,0.2]$ meV [Fig. 5(d)]. Upon cooling into the spin-frozen state at $T=0.05$ K, the $\bm{Q}$-dependence of the elastic scattering changes dramatically, as shown in Fig. 5(e). The resulting pattern is reminiscent of the magnetic diffuse scattering observed in CSI materials with strong dipolar interactions such as $\text{Ho}_{2}\text{Ti}_{2}\text{O}_{7}$[46, 47] and $\text{Dy}_{2}\text{Ti}_{2}\text{O}_{7}$[48, 49], but deviates even more from the ideal pinch-point structure. Instead, the scattering becomes concentrated near reciprocal space positions associated with a $\bm{q}=(100)$ propagation vector, suggesting a tendency towards incipient $\bm{q}=(100)$ order as the system undergoes freezing. The $\bm{Q}$-dependence of the gapped excitation observed at $T=0.05$ K [Fig. 5(a)] is revealed in Fig. 5(g), showing a starfish-shaped distribution reminiscent of a smeared pinch-point pattern, similar to inelastic magnetic features previously observed in several QSI candidates [23, 50, 22, 27]. A similar $\bm{Q}$-dependence is observed for the quasielastic scattering at $T=0.5$ K over the same energy range $\hbar\omega\in[0.2,0.6]$ meV [Fig. 5(f)], although the features appear more diffuse. For higher energy transfers $\hbar\omega\geq 0.6$ meV [Fig. 5(h, j)], no significant $\bm{Q}$-dependence is observed. Additional data are presented in Figs. A7,  A9, and A10.
@@ IV Discussion
The second mechanism assumes that the spin-freezing transition is driven by the spatial variance of the transverse field, $\delta h$, exceeding a critical threshold $\delta h_{c}$ [29, 30]. In this regime, certain ring-exchange terms become significantly enhanced relative to others, favoring the condensation of vison excitations. We interpret the observed suppression of quasielastic scattering below $T_{f}$ as a signature of vison condensation, which suppresses the low-energy quantum fluctuations characteristic of a quantum spin ice (QSI). Theoretically, vison condensation has also been predicted to induce $(100)$ magnetic order [6], which indeed appears incipient in the frozen state [Fig. 5(e)].
@@
Our results mark an important step toward realizing a U(1) QSL state, as they indicate that $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$ lies in close proximity to the ideal U(1) QSL in two key respects. First, in the temperature range above $T_{f}$ but below the onset of spin-ice correlations at $\sim$1 K, $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$ exhibits hallmarks of a QSL state: the characteristic wavevector dependence of the diffuse scattering [40] and the presence of two well-separated magnetic timescales [36]. Second, the spin-frozen state that develops below $T_{f}$ appears to emerge from the QSL via a confinement transition, in contrast to floating-zone–grown $\text{Pr}_{2}\text{Zr}_{2}\text{O}_{7}$ samples [21, 23, 22], which remain in a paramagnetic regime separated from the QSI phase by a first-order transition.
@@ Acknowledgements
This manuscript has been authored by UT-Battelle, LLC, under contract DE-AC05-00OR22725 with the U.S. Department of Energy (DOE). The U.S. government retains, and the publisher—by accepting the article for publication—acknowledges that the U.S. government retains a nonexclusive, paid-up, irrevocable, worldwide license to publish or reproduce the published form of this manuscript, or to allow others to do so, for U.S. government purposes. DOE will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan (https://www.energy.gov/doe-publicaccess-plan).
@@ Appendix A Model for the gapped excitation in spin-frozen state
As discussed in Section IV, the spin-frozen phase below $T_{f}\approx 0.15$ K hosts a gapped excitation at $\hbar\omega=0.23$ meV with a starfish-like $\bm{Q}$ dependence, reminiscent of inelastic features reported in other QSI candidates [23, 50, 22, 27]. In $\text{Pr}_{2}\text{Zr}_{2}\text{O}_{7}$, analogous excitations were attributed to CEF transitions between singlets of a split non-Kramers doublet, with the $\bm{Q}$ dependence shaped by finite nearest-neighbor correlations $J_{1}$ [22, 30]. Within a mean-field framework [23, 24], the elastic and inelastic spectral weights scale as $|\cos\theta_{i}|^{2}$ and $|\sin\theta_{i}|^{2}$, respectively, where $\theta_{i}$ is the angle between the pseudospin (defined in Eq. 1) and the local $\hat{\bm{z}}$ axis. Our finding that the inelastic channel carries only $\sim$40% of the total weight indicates a moderate deviation of the pseudospins from the Ising axis and suggests that the system retains features of the paramagnetic phase in the phase diagram [Fig. 1].
To characterize further the behavior of these CEF excitations, we follow Ref. [24] and introduce a mean-field model in which disorder in Ising pyrochlores with a non-Kramers ground-state doublet is represented by a site-dependent transverse field $h_{i}$ acting on each Pr3+ ion [29, 30]. The system is described by the Hamiltonian given by Eq. 4.
When $h_{i}$ is finite, the pseudo-spin $\sigma_{i}$ at site $i$ is no longer aligned strictly along the local $\langle 111\rangle$ $\hat{\bm{z}}$ direction. Following Refs. [23, 24], we consider a local mean-field Hamiltonian for site $i$,
@@
where $a_{1}$, $b_{1}$, $a_{2}$, and $b_{2}$ are normalized coefficients, satisfying $|a_{1}|^{2}+|b_{1}|^{2}=1$ and $|a_{2}|^{2}+|b_{2}|^{2}=1$. Compared to the convention adopted in Ref. [23], where a local spherical angle $\theta_{i}$ is defined between the pseudospin and the $\hat{\bm{z}}$ direction, the parameters used here are related by: $a_{1}\simeq\sin(\theta_{i}/2)$, $b_{1}\simeq\cos(\theta_{i}/2)$, $a_{2}\simeq\cos(\theta_{i}/2)$, and $b_{2}\simeq-\sin(\theta_{i}/2)$.
Since neutrons primarily couple to the $\hat{\bm{z}}-$component of the real spin for systems with non-Kramers ground state doublet [7], and the projected spin operator $J_{z}$ within the non-Kramers doublet is proportional to the pseudo-spin operator $\sigma_{z}$ [23], the neutron scattering intensities in the elastic and inelastic channels can be expressed as:
@@ Appendix B Subtraction of nuclear Schottky anomaly in heat capacity
We analyzed the nuclear Schottky anomaly in $C_{p}$ using two models: (1) the analytical model of Kimura et al. [21], with $\mu_{\mathrm{hyp}}^{\mathrm{Pr}}=1.9~\mathrm{\upmu_{B}}$ and no scaling factor, as in Ref. [21, 35]; and (2) a statistical model (Gopal model) [43] treating the system as a multilevel scheme with evenly spaced energy levels, using a level spacing of $\epsilon_{r}=0.011$ meV. Both models successfully reproduce the broad maximum in $C_{p}$ near 0.1 K. We note that our heat capacity data and the parameters obtained from model (1) are very similar to those of $\text{Pr}_{2}\text{Zr}_{2}\text{O}_{7}$ sample 1 reported in Ref. [44]. The temperature dependence of the magnetic entropy $S_{m}$, shown in Fig. 4(ii), was determined by subtracting both the Gopal-model nuclear contribution and the lattice contribution (estimated from nonmagnetic La2Sn2O7) from the measured $C_{p}$, and integrating the resulting magnetic contribution $C_{m}/T$. We emphasize that subtraction using either the Kimura or the Gopal model yields essentially identical results for the magnetic entropy.
@@ Appendix C Additional information for $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$ samples
Magnetization versus field measurements at $T=2$ K on single-crystal $\text{Pr}_{2}\text{Sn}_{2}\text{O}_{7}$  samples with masses of approximately 5 mg are shown in Fig. A4(c). The induced magnetic moments $M$ at the maximum applied field $\mu_{0}H=7$ T are measured to be $1.64~\mu_{\mathrm{B}}/\mathrm{Pr^{3+}}$, $1.33~\mu_{\mathrm{B}}/\mathrm{Pr^{3+}}$, and $1.19~\mu_{\mathrm{B}}/\mathrm{Pr^{3+}}$ for fields applied along the $[100]$, $[111]$, and $[110]$ directions, respectively. For comparison, the expected saturated moments in a pyrochlore lattice with Ising anisotropy along the local $\langle 111\rangle$ directions and effective ferromagnetic (FM) coupling between nearest neighbors are $g_{J}J/\sqrt{3}$ for the $[100]$ direction, $g_{J}J/2$ for $[111]$, and $g_{J}J/\sqrt{6}$ for $[110]$ [51]. Both the observed ordering of the induced moments (from largest to smallest) and their relative ratios are consistent with the predictions for a magnetic pyrochlore with easy-axis FM anisotropy. Quantitative deviations may arise from slight misalignments during sample mounting, mass uncertainty due to residual impurity phases, or the fact that full saturation may not have been achieved at $\mu_{0}H=7$ T. The temperature dependence of the dc magnetic susceptibility was measured with the magnetic field applied along the $[100]$, $[111]$, and $[110]$ crystallographic directions using the same single-crystal samples. To obtain a representative bulk behavior, we compute a powder-averaged dc susceptibility for a cubic system using the expression $\chi_{\mathrm{powder}}^{\mathrm{DC}}=(3\chi_{100}^{\mathrm{DC}}+6\chi_{110}^{\mathrm{DC}}+4\chi_{111}^{\mathrm{DC}})/13$ [52], and plot the results in Fig. A4(d,e). This powder-averaged dc susceptibility is also used to normalize the real part of the ac susceptibility $\chi^{\prime}$ presented in Fig. 5(a). A Curie–Weiss fit to the high-temperature regime yields a Curie–Weiss temperature of $\theta_{\mathrm{CW}}=-0.23(2)$ K and an effective moment of $\mu_{\mathrm{eff}}=2.78(1)~\mu_{\mathrm{B}}/\mathrm{Pr^{3+}}$, in good agreement with the powder measurements reported in Ref. [35].
To extract the “slow” timescale of spin dynamics [36] from the real part $\chi^{\prime}$ of the ac susceptibility, we employ a damped harmonic oscillator model [42, 35]: